Shift summary for day shift, 18/Dec/99

AG (ag@fnal.gov)
Sat, 18 Dec 1999 17:04:12 -0500 (EST)

Shift Summary for day shift, 18/Dec/99

On Shift: MA, AG

Hours of beam available at all intensities: 0.0h
Hours of beam available at nominal intensity: 0.0h
Hours of E799 data (special runs not included): 0.0h
Hours of special runs (laser scans, muon runs...): 0.0h

Accounting of lost time:

o accelerator is down.

========================================================================

Runs ended this shift:

Nothing useful.

Tapes brought to FCC this shift:
(including tapes for on-line split)

none.

========================================================================

Accesses during shift:

Interlocks are dropped since the last evening.

========================================================================

Overview of shift:

Recovering from power failures.

========================================================================

Description of Shift:

Taiji visited KTeV and explained the situation with power failures. Looks
like things have stabilized and there is no need for any other power
outages. We decide to turn on readout crates and take care of other
systems. Doug comes in and starts pumping the vacuum tank, so that the
vacuum is ~ 2.E-6 by 11:00. The valve on one DP isn't working so we're
running with one DP,which is OK.

At 1200 CsI lv/pip-line readout established by BobT. He also installed
fake spill signal, and removed accelerator spill signal. Large Bryair unit
is still not working properly - Thermflo will send Larry to come out
today. Small Bryair unit is keeping humidity under control.

13:15 Analysis magnet power supplies locked out in NS7 service building.

L1 diagnostics shows that V'-east RF strobe is bad (Grp5, src4)
-> Replaced two 622 coincidence modules and moved SA3 back to group 1
module. L1 diagnostics all pass.

Larry from Thermflo is here and continue work on large Bryair
unit. John Kent and Rich Schimdt are also working on it, there
is some strangeness in communication between PLC and the Bryair.

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift

CsI clock fanout has problems in RF patches:

5TE5, 11TE4, 10TW(1,2), 11TW5.

P. Shanahan will address these problems at ~18:00.